find class boundaries, midpoint, and width for the class

18.8 - 22.6


1
Expert's answer
2022-02-01T16:49:00-0500

Lower Boundary: 18.8-0.05=18.75.

Upper Boundary: 22.6+0.05=22.65.


Midpoint: "\frac{18.8+22.6}{2}=20.7."


Width: 22.65-18.75=3.9.